OpenStudy (anonymous):

A manufacturer wants to design an open top box having square base and a surface area of 48 square inches. What dimensions will produce a box with maximum volume? The height of the box should be inches, and the base of the box should be × inches.

OpenStudy (anonymous):

s=48 s=x^2+4xh V=x^2h

OpenStudy (anonymous):

so 48=x^2+4xh

jimthompson5910 (jim_thompson5910):

it's probably easier to solve for h in 48=x^2+4xh

jimthompson5910 (jim_thompson5910):

then plug this into V=x^2h then derive to find the critical values, which will lead you to the max volume (with corresponding dimensions)

OpenStudy (anonymous):

so 48/(x^2+4x)=h

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

more like 48=x^2+4xh 48-x^2=4xh h = (48 - x^2)/(4x)

OpenStudy (anonymous):

ah ok my bad, been a long day lol

OpenStudy (anonymous):

V=x^2((48 - x^2)/(4x)) V'=12-(3 x^2)/4

OpenStudy (anonymous):

got it thanks

jimthompson5910 (jim_thompson5910):

V=x^2h V=x^2((48 - x^2)/(4x)) V = (1/4)*x(48-x^2) V = (x^3)/4 - 12x V ' = (3x^2)/4 - 12 0 = (3x^2)/4 - 12 solve for x

jimthompson5910 (jim_thompson5910):

oh ok, glad you did

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yw